Aug 14, 2012 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US+Canada
    • +1-303-248-0285 : International
    • then enter participant code: 3421244# (remember the "#").
    • or (and code without the #)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6337 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting(s)

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. Monitoring
    1. LED system on Mini BCAL
  4. FCAL
    1. Construction Status
  5. BCAL
    1. Activities at JLab (Module positioning, SiPM binning, LG Travellers)
    2. Work at Regina
  6. USM update
    1. SiPM Update (Javiera)
    2. Light guides (Iñaki)
  7. TDR update
    1. Instructions for the TDR (You must be logged into the wiki to access this page).
    2. We need to revise Chapter 4 Detector, Section 4.3 Calorimetry.
  8. Calibration of Calorimeters
  9. Simulations/Reconstruction
  10. Any other business


Attendees: John, Beni, David, Yi, elton, Tim, Eugene, Manuel, Dan (JLab); Hayk, Javiera, Iñaki, Will (USM); Zisis, Andrei (UofR); Matt (IU)l. Thanks to John for running the meeting.

  1. Announcements
  2. Review of Action Items
  3. Monitoring
    1. Update from Athens
      • All the components for the Bcal boards should be in by Monday
      • The boards will start to be stuffed the week of the 27.
      • They hope to have the first shipment to JLab in mid Sep. This might match production shipments of light guides from Chile
      • Fcal: Still waiting on review by Fernando on the monitoring boards.
    2. Installation of monitoring boards (Elton)
      • Elton and Tara Bogart continued to investigate the options for installing the Bcal monitoring boards after gluing the light guides
      • The new jig holder for the guides is sufficiently thin and narrow that the holder and boards fit between light guides
      • The LED boards, which had lost their LEDs, had a resistor soldered in its place. The resistors were about 50% smaller than the LEDs in width and height.
      • We glued the boards into place with the UV curing glue. One of the pockets had something in it. Seven of the nine resistors were visible in the pocket through the light guide. After trying to reglue one of the resistors into the end light guide, several resistors popped out of their pocket. Only five resistors remained in their pocket. We removed the boards and found that the UV curing glue had not cured, except in two locations. John said that we should check for expired glue and possibly degradation of light source.
      • We attempted to glue the string a second time. This time we shinned the light through the top of the light guide and in between the cover and guide. This seemed to work better. We will inspect the second gluing job later today. It was more difficult to see the resistors clearly in the pockets on this attempt likely due to glue accumulating in the pocket.
    3. Addendum - added to minutes on 8/15 (Elton and Tara)
      • See photos at BCAL Monitoring
      • We checked the expiration date on the glue. It is Feb 2013
      • The light source has 50 hours of operation, still more to go
      • We checked that the glue does cure if it is directly illuminated for the nominal time
      • We illuminated one card for twice the amount of time on the end location (one with best access). The glue did cure, although the area was still wet. It seems that we can cure the glue, but care must be taken to illuminate for a longer period of time and/or try to illuminate from the inside. This will require further experimentation to develop a robust procedure.
    4. Discussion
      • Zisis: The schedule for the light guides and monitoring boards are starting to fall into line and we might have the boards at the time we glue the light guides
      • Tim: We will continue to pursue gluing the monitoring string after guides are attached as a backup plan.
  4. FCAL
    1. Bases are moving to completion. Almost at a point, where company has delivered all units. However, an accounting needs to take place including units that have been returned for rework
    2. Cable map. Map has been rearranged crate locations. Fernando has received the revised list, but has not acknowledged it yet. BCAL Latest version uploaded to portal
      • Labeling supplies coming in this week and labeling can begin.
    3. Support for dark room. Fitting all Al plates. John is happy with the bottom half. Top half will be checked next week.
    4. Work on darkroom design is ongoing. Leckey and Fry have developed concept and now moving to design with full 3D CAD model. Would like to complete the design so that any additional drilling or tapping can be done before shipping.
      • To finalize the design we need to pin down the mounting system. Chuck has sent some pptx files, but Tim said we had sent drawings to Athens for review. Athens would like square sheets (except for the beamhole). We will update drawings and send them around for comments. John Fry can develop the concept for attaching/hanging.
    5. Installation schedule is being worked through. This will be checked against the master schedule and then set data for shipping of Fcal frame.
    6. Activities at JLab (Module positioning, SiPM binning, LG Travellers)
      • Tegan found a solution on the placement of modules inside barrel
      • Shaun worked on the assignment of SiPMs to electronic boards
      • They are now back a Regina and will start taking scans of laser intensity at several bias settings and check resolution
    7. Work at Regina
      • Andrei has finished taking the cosmic runs over a several bias settings
  5. USM update
    1. SiPM Update (Javiera)
      • Elton: when will the last shipment from Hamamatsu be sent? They are still working out some custom issues. When resolved, they will have the shipment sent, but it is not yet scheduled.
      • Elton: The second shipment to JLab arrived without a hitch. Future shipments will follow the same procedure.
    2. Light guides (Iñaki)
      • Final calibrations have been completed. Pieces are within tolerance, except on the length dimension (80 mm).

Question: Can these be used? Iñaki will send tables with data. Elton and Tim will review for acceptability

  1. TDR update
    1. Instructions for the TDR (You must be logged into the wiki to access this page).
      • Zisis will continue working on the TDR this month.
      • Matt: Wanted to check that Bcal and Fcal sections are in different files to ease parallel editing. This seems to be the case.
  2. Calibration of Calorimeters
  3. Simulations/Reconstruction
    1. Andrei: Simulation runs are almost complete. A final check between GEANT3 and FLUKA will be done to verify agreement.
    2. David: What are the parameters for the resolution? Andrei: The runs are taken at different 1) energies 2) polar angle 3) distance from center of cell (phi) 4) depth of cell. He envisions having tables with fine and smooth granularity in the first three variables at fixed depths inside the module.
    3. David: Some issues were revealed when his studies were shown at an earlier meeting. These were the result of a 45 mV threshold that were included in the mcsmear. The threshold should be much lower for the ADC signal. The thresholds were affecting the resolution parameters. The program is being modified to have two thresholds, one for the ADC and one for the TDC. The ADC threshold is effectively 5-10 times smaller.
  4. Any other business